OpenStudy (anonymous):

Kellogg's sells cereal in a box that has dimensions 8 inches in length, by 2 inches in width, by 9 inches in height If the company decides to sell a larger cereal box that is similar to their smaller box, what will be the height of the larger box if it's length is now 10 inches? A ) 2 1/2 in B ) 11 1/4 C ) 11.4 in D ) 17 in

OpenStudy (anonymous):

Help me please

OpenStudy (anonymous):

Both boxes are similar hence their lengths widths and heights will be in proportion therefore, \[\huge \frac{8}{ 10}= \frac{9}{x} \rightarrow x= \frac{ 10 \times 9}{8}\] \[\huge \rightarrow x= \frac{ 5 \times 9}{4} = \frac{45}{4} \] \[\huge 11 \frac{1}{4} \] hence option B ) 11 1/4 is the correct answer. @Nicole_xoxo

OpenStudy (anonymous):

Thank u